A car moves at a constant acceleration and crosses a path of 400m. What is the acceleration if the car takes 10 seconds to cross the third 100m path?

1
Expert's answer
2021-09-13T11:02:12-0400

"\text{let for point } P \text{ middle of the 400m path}"

"x_0= 0;v_0>0"

"s(t) = v_0t +\frac{at^2}{2}"

"s(10)=100"

"s(10)=10v_0+50a"

"v_0+5a-10 = 0\ (1)"

"s_1=\frac{v_1^2-v_0^2}{2a}"

"\text{for }s_1=200(\text{ the car has reached the end of the 400m path}):"

"v_1 = 0"

"200=\frac{-v_0^2}{2a}"


"a= -\frac{v_0^2}{400}"

"\text{Substitute a into expression (1)}"

"v_0-\frac{v_0^2}{80}-10 = 0"

"v_0^2-80v_0+800 = 0"

"v_0= 11.72\frac{m}{s}; a = -\frac{v_0^2}{400}= -0.34\frac{m}{s^2}"

"v_0'= 68.28\frac{m}{s}; a = -\frac{v_0^2}{400}= -11.6\frac{m}{s^2}"


"\text{Answer: }-0.34\frac{m}{s^2} \ or\ -11.6\frac{m}{s^2}"